The massless spring of a spring gun has a force constant k=12N/cm. When the gun is aimed vertically, a 15-g projectile is shot to a height of 5.0 m above the end of the expanded spring. How much was the spring compressed initially?


1
Expert's answer
2021-03-14T19:15:44-0400

We can find the initial compression of the spring from the law of conservation of energy:


"\dfrac{1}{2}dx^2=mgh,""d=\sqrt{\dfrac{2mgh}{k}},""d=\sqrt{\dfrac{2\cdot0.015\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot5\ m}{12\ \dfrac{N}{cm}\cdot\dfrac{100\ cm}{1\ m}}}=0.035\ m=3.5\ cm."